# The Architecture and Evaluation of Bayesian Neural Networks

### v1

@@ 2.5 Out-of-Distribution Prediction
While it is not surprising that the accuracy and the quality of uncertainty quantification of any model decreases under a distribution shift, reliable uncertainty estimates that are robust to the out-of-distribution (OOD) data become exceptionally important in safety-critical applications. The challenge is especially intricate since better accuracy and lower calibration error of a certain model on the in-domain data do not imply better accuracy and lower calibration error in the OOD settings [Ovadia et al., 2019]. Here, we wish to validate the models’ predictive abilities when the test data points come from previously unseen regions of data space. The kind of out-of-distribution data we consider could be described as ’complement-distributions’, such data arises in open-set recognition or could be the result of adversary [Farquhar and Gal, 2022]. Note that in Section C.1 we introduce a much milder example with ’related-distributions’ test data. We split the training data used in Sections 2.3 and 2.4 into the train and test data covering complement regions of the function. Specifically, $\bm{\mathcal{D}}=\bm{\mathcal{D}}_{c}\sqcup\tilde{\bm{\mathcal{D}}}_{c}$, the observed data $\bm{\mathcal{D}}_{c}$ consists of $N=370$, the new data $\tilde{\bm{\mathcal{D}}}_{c}$ consists of $\tilde{N}=130$ and the observed and the new data are disjoint (see Figure 4):
@@
Strictly speaking, we do not expect any model to be robust to such an extreme case and, mainly, want to asses and better understand the quality of the uncertainty estimates. In this experiment, we are hoping that the relationship between the distributions of the observed and the new data makes this challenge somewhat tractable. On Figure 4(a) we illustrate the metrics for $D_{1}=20,200,1000$ and $2000$ hidden units; Figure 4(b) compares non-OOD and OOD predictions obtained by the BNNs with ReLU activation and $D_{1}=200$. The poor performance of the mfVIS, especially for wider networks is not surprising, however, we notice that HMCS suffers from much higher RMSE than mfVIR and HMCR for wide networks. And while HMCR has a lower RMSE than any model trained with mean-field VI, the ability of HMC to capture the uncertainty deteriorates and it becomes overconfident. Whilst HMCR200 and mfVIR200 do not show any of the expected increase in the uncertainty, on certain regions both methods are able to provide accurate predictive mean (see Figure 4(b), the RHS region of the function, where $x>1.5$). Finally, as the width of the network increases, mfVIR outperforms all of the models.
(a) Out-of-sample predictive performance of all the models compared.
(b) Within-the sample and out-of-sample predictions and uncertainty estimates of BNNs with $D_{1}=200$ and ReLU activation.
@@ 3.1 Predictive Methods for Model Assessment
When considering synthetic datasets, we can choose a desired metric and sample any number of data points, so that evaluation of the model’s performance becomes trivial. For example, in Section 2.5 we have specifically created an extreme case when the training data $\bm{\mathcal{D}}_{c}$ and the new data $\tilde{\bm{\mathcal{D}}}_{c}$ were covering disjoint regions of the true function. In reality, the new previously unseen data is not available, and one can only estimate the expected out-of-sample predictive performance. Suppose that we only observe $\bm{\mathcal{D}}$, the unseen observations $\tilde{\bm{\mathcal{D}}}$ are generated by $p_{t}(\tilde{\bm{\mathcal{D}}})$, and we wish to be able to access the generalization ability of the model without having access to the test data. To keep the notation simple we omit the dependency on $\bm{x}$ and $\tilde{\bm{x}}$ when writing down the posteriors in this section. Given a new data point $\tilde{y}_{n}$, the log score $\log p(\tilde{y}_{n}|\bm{\mathcal{D}})$ is one of the most common utility functions used in measuring the quality of the predictive distribution. The log score benefits from being a local and proper scoring rule [Vehtari and Ojanen, 2012]. Then, the expected log pointwise predictive density for a new dataset serves as a measure of the predictive accuracy of a given model:
@@
where $p(\tilde{y}_{n}|\bm{\mathcal{D}})$ is model’s posterior predictive distribution. In the absence of $\tilde{\bm{\mathcal{D}}}$, one might obtain an estimate of the expected log pointwise predictive density by re-using the observed $\bm{\mathcal{D}}$. Here, we review the approach that employs leave-one-out cross-validation (LOO-CV), which can be seen as a natural framework for accessing the model’s predictive performance [Vehtari et al., 2016].
Willing to obtain the Bayesian leave-one-out cross-validation (LOO-CV) estimate of the expected utility $\widehat{\text{elpd}}_{\text{loo}}$ and avoid re-fitting the model $N$ times one could use importance sampling. However, the classical importance weights would have a large variance and the obtained estimates would be noisy. Recently, the problem was solved with the Pareto smoothed importance sampling (PSIS) which allows evaluating the LOO-CV expected utility in a reliable yet efficient way [Vehtari et al., 2016]:
@@
where $r_{i}^{s}$ are the smoothed importance weights which benefit from smaller variance than the classical weights. We refer to the individual logarithms in the sum as to $\widehat{\text{elpd}}_{\text{loo},n}$. The advantage of PSIS is that the estimated shape parameter of the Pareto distribution provides a diagnostic of the reliability of the resulting expected utility. Although the methods of model selection which reuse the data can be vulnerable to overfitting when the size of the dataset is too small and/or the data is sparse, it is (relatively) safe to use cross-validation to compare a small number of models and given a large enough dataset [Vetari et al., 2019]. In Section B.1, we implement $\widehat{\text{elpd}}_{\text{loo}}$ in the empirical experiment, where we additionally consider posterior predictive checks (PPC) and an alternative to the LOO-CV approach of estimating the expected log pointwise utility.
@@ 3.2 Alternatives to Classical Bayesian Model Averaging
Let =$\mathcal{M}=\{M_{1},\ldots,M_{K}\}$ be a collection of models and denote the parameters of each of the $M_{k}$ as $\bm{\theta}_{k}$. The assumptions one has on the prediction task and on $\mathcal{M}$ with respect to the true data-generating process, can be categorized into three scenarios: $\mathcal{M}$-closed, $\mathcal{M}$-open and $\mathcal{M}$-complete. If $M_{k}\in\mathcal{M}$ for some $k$ recovers the true data generating process then we are in the $\mathcal{M}$-closed case. The task is $\mathcal{M}$-complete if there exists a true model but it is not included in $\mathcal{M}$ (e.g. for computational reasons). Finally, we are in the $\mathcal{M}$-open scenario when the true model is not in $\mathcal{M}$ and the data generating mechanism cannot be conceptually formalized to provide an explicit model [Vehtari and Ojanen, 2012]. The Bayesian framework allows to define the probabilities over the model space and the $\mathcal{M}$-closed case, the classical Bayesian Model Averaging would be able to give optimal performance. The BMA solution provides an averaged predictive posterior as [Hoeting et al., 1999]
@@
However, in the $\mathcal{M}$-open and $\mathcal{M}$-complete prediction tasks, the BMA is not appropriate as it gives a strong preference to a single model and so assumes that this particular model is the true one. Now, if we replace the weights $p(M_{k}|\bm{\mathcal{D}})$ with the products of Bayesian LOO-CV densities $\prod_{n=1}^{N}p(y_{n}\mid x_{n},\bm{\mathcal{D}}_{-n},M_{k})$, we arrive at pseudo-Bayesian model averaging (pseudo-BMA). In other words, the weights $w_{k}$ of pseudo-BMA are proportional to the estimated log pointwise predictive density $\exp(\widehat{\text{elpd}}_{\text{loo}}^{k})$ introduced in Section 3.1. One could further correct each $\widehat{\text{elpd}}_{\text{loo}}$ estimate of Equation 2 by the standard errors and obtain
@@
where a PSIS estimate of the predictive LOO-CV density is used and $r_{i}^{s}$ are the smoothed (truncated) importance weights.
Finally, we recall that deep ensembles of classical non-Bayesian NNs [Lakshminarayanan et al., 2017] behave similarly to Bayesian model averages, and both lead to solutions strongly favouring one single model [Wilson and Izmailov, 2020]. In contrast, the ensembles of BNN posteriors in the Equation 3 with $p(M_{k}\mid\bm{\mathcal{D}})=K^{-1}$ can be seen as a trivial case of BMA which combines models and does not give preference to a single solution. Alternatively, when implementing variational inference and combining BNNs, the analogy can be drawn with the simplified version of the adaptive variational Bayes, which combines variational posteriors with certain weights and under certain conditions, attains optimal contraction rates [Ohn and Lin, 2024].
@@ 3.3 Ensembles and Averages
We compare three model averaging methodologies: deep ensembles of Bayesian neural networks, stacking and pseudo-BMA based on PSIS-LOO [Yao et al., 2018]. We do not consider the Bayesian Bootstrap (BB) [Rubin, 1981] motivated by the recent observation that in the settings of modern neural networks deep ensembles of non-Bayesian NNs and BB are equivalent, and both are often misspecified [Wu and A Williamson, 2024]. Combining several estimates of BNNs can be effective not only when predictions are coming from different models, but also when dealing with several predictions obtained by the same model [Ohn and Lin, 2024]. This is of particular use for multimodal posteriors arising in BNNs where different modes could be explored by random initializations [Yao et al., 2018]. Additionally recall, that the ELBO, the objective of variational inference, is a non-convex function so that the optimum is only local and depends on the starting point. We note that combining models trained with HMC and VI would be meaningless for several reasons. First of all, training a set of HMC models becomes rather expensive: for instance, training the HMCR20 once takes the same amount of time as 35 trainings of mfVIR20. Second, the estimates of the log pointwise predictive densities (provided in Section B.1) for HMC and VI have different scales and are not easily compared, in this case, the result of averaging HMC and VI would be equivalent to classical BMA. The reader, nevertheless, interested in ensembles and averages of HMC runs is referred to Section C.2 where we show that in this particular example, neither of the methods promoted exploring the multiple modes of the posterior predictive distribution nor did they help improve uncertainty quantification. Additionally Section C.3 provides results of ensembling and averaging the deeper networks and Section C.1 tests the performance of averaging methods in less extreme examples of OOD data by introducing a much milder example with the so-called ’related-distributions’ test data Farquhar and Gal [2022].
@@
Now consider the mfVIR20 model and the ’complement-distributions’ data of Section 2.5. We choose 10 random initialization points, obtain 10 posterior predictive distributions and compute estimated expected log pointwise predictive densities. We then construct ensemble, pseudo-BMA and stacking approximations, the results are illustrated Figure 5). The ensembling and stacking provide subtle results and are superior to pseudo-BMA which has worse accuracy and fails to capture any uncertainty..
@@
Figure 5: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20. The pseudo-BMA is worse than DE and stacking which are very similar to each other.
@@
General summary. We observe that, similarly to BMA, the pseudo-BMA is not preferable in $\mathcal{M}$-open and $\mathcal{M}$-complete settings. Namely, in the experiments the pseudo-BMA was confirmed to be inferior to stacking and ensembles of BNNs both, in terms of the predictive accuracy and the uncertainty quantification. Stacking and ensembles of BNNs performed comparable to each other and provided an improvement, which is especially significant in terms of the uncertainty quantification.

### v2

@@ 2.5 Out-of-Distribution Prediction
While it is not surprising that the accuracy and the quality of uncertainty quantification of any model decreases under a distribution shift, reliable uncertainty estimates that are robust to the out-of-distribution (OOD) data become exceptionally important in safety-critical applications. The challenge is especially intricate since better accuracy and lower calibration error of a certain model on the in-domain data do not imply better accuracy and lower calibration error in the OOD settings [Ovadia et al., 2019]. Here, we wish to validate the models’ predictive abilities when the test data points come from previously unseen regions of data space. The kind of out-of-distribution data we consider could be described as ’complement-distributions’, such data arises in open-set recognition or could be the result of an adversary [Farquhar and Gal, 2022]. Note that in Section 3.3 as well as in the supplementary, we consider a much milder example with ’related-distributions’ data. We split the training data used in Sections 2.3 and 2.4 into the train and test data covering complement regions of the function.
Specifically, $\bm{\mathcal{D}}=\bm{\mathcal{D}}_{c}\sqcup\tilde{\bm{\mathcal{D}}}_{c}$, the observed data $\bm{\mathcal{D}}_{c}$ consists of $N=370$, the new data $\tilde{\bm{\mathcal{D}}}_{c}$ consists of $\tilde{N}=130$ and the observed and the new data are disjoint (see Figure 4):
@@
Strictly speaking, we do not expect any model to be robust to such an extreme case and, mainly, want to asses and better understand the quality of the uncertainty estimates. In this experiment, we are hoping that the relationship between the distributions of the observed and the new data makes this challenge somewhat tractable. On Figure 4(a) we illustrate the metrics for $D_{1}=20,200,1000$ and $2000$ hidden units; Figure 4(b) compares non-OOD and OOD predictions obtained by the BNNs with ReLU activation, Gaussian priors and $D_{1}=200$. The poor performance of the mfVIS, especially for wider networks, is not surprising, however, we notice that for wide networks HMCS with Gaussian priors suffers from much higher RMSE than HMCS with Student-t priors and mfVIR and HMCR with either choice of priors. And while HMCR has a lower RMSE than any model trained with mean-field VI, the ability of HMC to capture the uncertainty deteriorates, and it becomes overconfident. Whilst HMCR200 and mfVIR200 do not show any of the expected increase in the uncertainty, on certain regions both methods are able to provide accurate predictive mean (see Figure 4(b) for examples with Gaussian priors, the right-hand side region of the function, where $x>1.5$). Finally, as the width of the network increases, mfVIR outperforms all of the methods.
(a) Out-of-sample metrics of all the methods compared.
(b) Within-the sample (left) and out-of-sample (right) predictions and uncertainty estimates for $D_{1}=200$, ReLU activation and Gaussian priors.
@@ 3.1 Predictive Methods for Model Assessment
When considering synthetic datasets, we can choose a desired metric and sample any number of data points, so that evaluation of the model’s performance becomes trivial. For example, in Section 2.5 we have specifically created an extreme case when the training data $\bm{\mathcal{D}}_{c}$ and the new data $\tilde{\bm{\mathcal{D}}}_{c}$ were covering disjoint regions of the true function. In reality, the new previously unseen data is not available, and one can only estimate the expected out-of-sample predictive performance. Suppose that we only observe $\bm{\mathcal{D}}$, the unseen observations $\tilde{\bm{\mathcal{D}}}$ are generated by $p_{t}(\tilde{\bm{\mathcal{D}}})$, and we wish to be able to assess the generalization ability of the model without having access to the test data. To keep the notation simple, we omit the dependency on $\bm{x}$ and $\tilde{\bm{x}}$ when writing down the posteriors in this section. Given a new data point $\tilde{y}_{n}$, the log score $\log p(\tilde{y}_{n}|\bm{\mathcal{D}})$ is one of the most common utility functions used in measuring the quality of the predictive distribution. The log score benefits from being a local and proper scoring rule [Vehtari and Ojanen, 2012]. Then, the expected log pointwise predictive density for a new dataset serves as a measure of the predictive accuracy of a given model:
@@
where $p(\tilde{y}_{n}|\bm{\mathcal{D}})$ is model’s posterior predictive distribution. In the absence of $\tilde{\bm{\mathcal{D}}}$, one might obtain an estimate of the expected log pointwise predictive density by re-using the observed $\bm{\mathcal{D}}$. Here, we review the approach that employs leave-one-out cross-validation (LOO-CV), which can be seen as a natural framework for assessing the model’s predictive performance [Vehtari et al., 2016].
To obtain the Bayesian leave-one-out cross-validation (LOO-CV) estimate of the expected utility $\widehat{\text{elpd}}_{\text{loo}}$ and avoid re-fitting the model $N$ times, one could use importance sampling. However, the classical importance weights would have a large variance, and the obtained estimates would be noisy. Recently, the problem was solved with Pareto smoothed importance sampling (PSIS), which allows evaluating the LOO-CV expected utility in a reliable yet efficient way [Vehtari et al., 2016]:
@@
where $r_{i}^{s}$ are the smoothed importance weights, which benefit from smaller variance than the classical weights. We refer to the individual logarithms in the sum as $\widehat{\text{elpd}}_{\text{loo},n}$. The advantage of PSIS is that the estimated shape parameter of the Pareto distribution provides a diagnostic of the reliability of the resulting expected utility. Although methods of model selection which reuse the data can be vulnerable to overfitting when the size of the dataset is too small and/or the data is sparse, it is (relatively) safe to use cross-validation to compare a small number of models and given a large enough dataset [Vetari et al., 2019]. In Appendix B, we implement $\widehat{\text{elpd}}_{\text{loo}}$ in the empirical experiment, where we additionally consider posterior predictive checks (PPC) and an alternative to the LOO-CV approach of estimating the expected log pointwise utility.
@@ 3.2 Alternatives to Classical Bayesian Model Averaging
Let =$\mathcal{M}=\{M_{1},\ldots,M_{K}\}$ be a collection of models and denote the parameters of each of the $M_{k}$ as $\bm{\theta}_{k}$. The assumptions one has on the prediction task and on $\mathcal{M}$ with respect to the true data-generating process can be categorized into three scenarios: $\mathcal{M}$-closed, $\mathcal{M}$-open and $\mathcal{M}$-complete. If $M_{k}\in\mathcal{M}$ for some $k$ recovers the true data generating process, then we are in the $\mathcal{M}$-closed case. The task is $\mathcal{M}$-complete if there exists a true model but it is not included in $\mathcal{M}$ (e.g. for computational reasons). Finally, we are in the $\mathcal{M}$-open scenario when the true model is not in $\mathcal{M}$ and the data-generating mechanism cannot be conceptually formalized to provide an explicit model [Vehtari and Ojanen, 2012]. The Bayesian framework allows to define the probabilities over the model space, and for the $\mathcal{M}$-closed case, classical Bayesian Model Averaging (BMA) would give optimal performance. The BMA solution provides an averaged predictive posterior as [Hoeting et al., 1999]
@@
However, in the $\mathcal{M}$-open and $\mathcal{M}$-complete prediction tasks, BMA is not appropriate as it gives a strong preference to a single model and so assumes that this particular model is the true one. Now, if we replace the weights $p(M_{k}|\bm{\mathcal{D}})$ with the products of Bayesian LOO-CV densities $\prod_{n=1}^{N}p(y_{n}\mid x_{n},\bm{\mathcal{D}}_{-n},M_{k})$, we arrive at pseudo-Bayesian model averaging (pseudo-BMA). In other words, the weights $w_{k}$ of pseudo-BMA are proportional to the estimated log pointwise predictive density $\exp(\widehat{\text{elpd}}_{\text{loo}}^{k})$ introduced in Section 3.1. One could further correct each $\widehat{\text{elpd}}_{\text{loo}}$ estimate of Equation 2 by the standard errors and obtain
@@
where a PSIS estimate of the predictive LOO-CV density is used, and $r_{i}^{s}$ are the smoothed (truncated) importance weights.
Finally, we recall that deep ensembles of classical non-Bayesian NNs [Lakshminarayanan et al., 2017] behave similarly to Bayesian model averages, and both lead to solutions strongly favouring one single model [Wilson and Izmailov, 2020]. In contrast, the ensembles of BNN posteriors in Equation 3 with $p(M_{k}\mid\bm{\mathcal{D}})=K^{-1}$ can be seen as a trivial case of BMA, which combines models and does not give preference to a single solution. Alternatively, when implementing variational inference and combining BNNs, the analogy can be drawn with the simplified version of adaptive variational Bayes, which combines variational posteriors with certain weights and, under certain conditions, attains optimal contraction rates [Ohn and Lin, 2024].
@@ 3.3 Ensembles and Averages
We compare three model averaging methodologies: deep ensembles of Bayesian neural networks, stacking and pseudo-BMA based on PSIS-LOO [Yao et al., 2018]. We do not consider the Bayesian Bootstrap (BB) [Rubin, 1981] motivated by the recent observation that in the settings of modern neural networks deep ensembles of non-Bayesian NNs and BB are equivalent, and both are often misspecified [Wu and A Williamson, 2024]. Combining several estimates of BNNs can be effective not only when predictions are coming from different models, but also when dealing with several predictions obtained by the same model [Ohn and Lin, 2024]. This is of particular use for multimodal posteriors arising in BNNs, where different modes could be explored by random initializations [Yao et al., 2018]. Additionally, recall that the ELBO, the objective of variational inference, is a non-convex function, so that the optimum is only local and depends on the starting point. We note that combining models trained with HMC and VI would be meaningless for several reasons. First of all, training a set of HMC models becomes rather expensive: for instance, training the HMCR20 once takes the same amount of time as 35 trainings of mfVIR20. Second, the estimates of the log pointwise predictive densities (provided in Appendix B) for HMC and VI have different scales and are not easily compared; in this case, the result of averaging HMC and VI would be equivalent to classical BMA.
@@
Now consider the mfVIR20 model with Gaussian priors and the ’complement-distributions’ data of Section 2.5. We choose 10 random initialization points, obtain 10 posterior predictive distributions and compute estimated expected log pointwise predictive densities. We then construct ensemble, pseudo-BMA and stacking approximations; the results are illustrated Figure 5. Ensembling and stacking are superior to pseudo-BMA, which has worse accuracy and fails to capture any uncertainty. Similar results for the mfVIR20 model with Student-t priors are provided in Appendix C.
@@
Figure 5: Predictions obtained by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with Gaussian priors in the ’complement-distributions’ (top) and ’related-distributions’ (bottom) OOD tasks. Pseudo-BMA is worse than the other methodologies, and stacking provides improvements over DE in uncertainty quantification.
Given the nature of the test data we use, the predictions as well as the $\widehat{\text{elpd}}_{\text{loo}}$ estimates may be unreliable. Thus, we consider a simpler data-generating mechanism in which test data comes from a slightly broader region; such a scenario could be called an OOD task with ’related-distributions’ [Farquhar and Gal, 2022]. Specifically, the data are generated as follows:
$\bm{x}$ $\sim\text{Unif}([0,1]),\quad\bm{y}=\sin(10\bm{x})\bm{x}^{2}+\bm{\epsilon},\quad\bm{\epsilon}\sim 0.05\text{N}(0,1).$
The data for training $\bm{\mathcal{D}}_{r}$ and the testing $\tilde{\bm{\mathcal{D}}}_{r}$ consist of $N=450$ and $\tilde{N}=50$ observations, respectively, where $\tilde{\bm{\mathcal{D}}}_{r}$ comes from the broader region than $\bm{\mathcal{D}}_{r}$, i.e. $(\min_{n=1\ldots N}(x_{n}),\max_{n=1\ldots N}(x_{n}))\subsetneq(\min_{n=1\ldots\tilde{N}}(\tilde{x}_{n}),\max_{n=1\ldots\tilde{N}}(\tilde{x}_{n}))$. For 10 posterior predictive distributions of mfVIR20 with Gaussian priors (results for Student-t priors are provided in Appendix C), we compare ensembling, pseudo-BMA and stacking in Figure 5 (similar results with having Student-t priors are presented in the supplementary, and in Section D.2 we additionally provide the results of ensembling and averaging in the deeper networks.). Whilst the total uncertainty estimates of pseudo-BMA are, somewhat, adequate, the model uncertainty is underestimated. Both stacking and deep ensembles lead to improved predictive performance and uncertainty quantification, with stacking showing some better gains compared to deep ensembles (see e.g. improved coverage of stacking on the right-hand side of Figure 5).
@@
General summary. We observe that, similar to BMA, the pseudo-BMA is not preferable in $\mathcal{M}$-open and $\mathcal{M}$-complete settings. Namely, in ’complement-distributions’ and ’related-distributions’ experiments, pseudo-BMA was confirmed to be inferior to stacking and ensembles of BNNs both in terms of the predictive accuracy and uncertainty quantification. Stacking and ensembles of BNNs performed comparable to each other and provided an improvement, with modest gains for stacking, which is especially significant in terms of uncertainty quantification in the OOD setting.
3.4 The Rocket Booster Simulation
We further evaluate the empirical performance of Bayesian model averaging techniques in BNNs by considering the data simulated by NASA when designing a novel rocket booster called Langley Glide-Back Booster (LGBB) [Gramacy and Lee, 2008, Rogers et al., 2003]. The data comprises 3167 observations, and the goal is to model the lift force as a function of the speed of the vehicle, the angle of attack $\alpha$, and the sideslip angle $\beta$. Speed ranges from Mach 0 to Mach 6, the angle $\alpha$ lies in between $-5$ and $30$ degrees and the sideslip angle $\beta$ takes values $0,0.5,1,2,3,4$.
We begin with the non-OOD scenario, consider a one-layer BNN with ReLU activation, Gaussian prior and $D_{1}=20,100$ or $200$ and train our model with mfVI on $80\%$ of the data. Figure 6 displays a slice of predicting lift, with Mach (speed) on the x-axis, $\alpha$ fixed to 25 and $\beta$ to 4; it compares predictions of a single mfVIR20 model to the results of combining 10 predictive distributions via deep ensembles, pseudo-BMA and stacking. Figures for slices of predicting lift obtained for models with $100$ and $200$ hidden units are provided in Section D.3. Figure 7 compares the metrics obtained by different approaches as the BNN gets wider. Stacking provides modest gains in accuracy, and while coverage is comparable across all methods, the slightly wider model uncertainty visible in the top row Figure 6 for stacking seems more reasonable given the observed data (crosses).
Figure 6: Slices of predicted lift obtained by a single model and by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with Gaussian priors in the non-OOD (top) and ’complement-distributions’ OOD (bottom) tasks.
Further, we construct the ’complement-distributions’ OOD scenario: the data taken for training consists of all datapoints but those whose sideslip angle $\beta$ equals 4, the data on which we then test the model consists of the remaining datapoints for which $\beta=4$. Similar to the non-OOD scenario, Figure 6 (bottom) compares slices of predicting lift obtained by a single mfVIR20 model to the predictions of deep ensembles, pseudo-BMA and stacking (figures for models with $100$ and $200$ hidden units are provided in Section D.3); this figure particularly highlights how a single model and pseudo-BMA can lead to incorrect, overconfident predictions in OOD settings. Unlike the non-OOD case, where performance is stable across the network’s width, it varies substantially for OOD (left versus right side of Figure 6), especially for pseudo-BMA and stacking, which show improvements in both accuracy and uncertainty as the width increases.
Figure 7: RMSE and CI obtained by a single mfVIR model, and by ensembling, stacking and pseudo-BMA when applied to mfVIR20 with Gaussian priors in the non-OOD (LHS) and ’complement-distributions’ OOD (RHS) tasks.
General summary. In the non-OOD scenario, stacking provides modest gains in both accuracy and uncertainty quantification, compared to deep ensembles and pseudo-BMA, with stable performance across different widths. Instead, for the ’complement-distributions’ OOD task, performance heavily depends on the network architecture. Deep ensembles show the most stability, with stacking and pseudo-BMA becoming competitive as the width increases.